OpenStudy (help_me_plz):

solve for x ----- 3x-5 divided by 7 = - x+2 divided by 4.

OpenStudy (saifoo.khan):

\[\frac{3x-5}{7}=\frac{-x+2}{4}\]This?

OpenStudy (help_me_plz):

yes

OpenStudy (help_me_plz):

the negative is before the second division question

OpenStudy (saifoo.khan):

Oh, that's the same thing.

OpenStudy (help_me_plz):

oh

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (saifoo.khan):

\[4(3x-5) = 7(-x+2)\]

OpenStudy (saifoo.khan):

Can you solve now?

OpenStudy (anonymous):

1.789473684

OpenStudy (help_me_plz):

would you like me to read off the multi choice answers?

OpenStudy (saifoo.khan):

Please try to solve it by yourself. It will help you Learn.

OpenStudy (help_me_plz):

all the multi choice answers are fractions also

OpenStudy (anonymous):

step1:cross multiply 4(3x-5)=7(-x+2) 12x-20=-7x+14 step2:Group x terms together and constants(i mean numbers) together. 12x+7x=14+20 19x=34 x=34/19 I hope this would help you solve similar problems